2 a) Find the particular solution for y - 2y + y = 6e b) Find y, for y + 3y - 36x² + 8e-> JT JT c)Find the general solut
Answer #1

Got 2 m) 4-2.4+ Auxillary earration Ł Complementary Farction qe=ciet + Caté &p= At?et y= zatet + At et q = zaek tua tet +-3t b) y+ 3y! =367tség Aurillary equation : D+30= SD=0-3 elp = AA3+ Bt²tet tot e 3t -3t -3t g= 3 At + 2Bt to toe - Bote - 3